the average cholesterol content of a certain canned goods is 215 milligram and the standard deviation is 15 milligram, answer the variable is normally distributed


1
Expert's answer
2021-05-13T06:55:32-0400

The complete question is :

The average cholesterol content of a certain canned goods is 215 milligrams, and the standard deviation is 15 milligrams. Assume the variable is normally distributed:

a.) If a canned goods is selected, what is the probability that the cholesterol content will be greater than 220 milligrams?

b.) If a sample of 25 canned goods is selected, what is the probability that the mean of the samle will be larger than 220 milligrams?

Answer:

a.) "P(X>220) = P(\dfrac{x- \mu}{\sigma}>\dfrac{220-215}{15})"


"= P(Z>0.333)"


"= 1-P(Z<0.333)"


"= 1-0.6304"


"= 0.369"


b.) We have "n = 25"


"P(\bar X >220) = 1-P(X<220)"


"= 1-P(Z<\dfrac{x-\mu}{\dfrac{\sigma}{\sqrt{n}}})"


"= 1-P(Z< \dfrac{220-215}{\dfrac{15}{\sqrt{20}}})"


"= 1-P(Z<1.49)"


"= 1-0.9318"

"= 0.068"